A certain airplane has two independent alternators to provide electrical power. The probability
that a given alternator will fail on a 1-hour flight is .02. What is the probability that (a) both will
fail? (b) Neither will fail? (c) One or the other will fail? Show all steps carefully.

Answer by stanbon(48568)

The probability
that a given alternator will fail on a 1-hour flight is .02. What is the probability that (a) both will
fail? (b) Neither will fail? (c) One or the other will fail? Show all steps carefully.
-----------
Prob(both fail) = 0.02^2 = 0.0004
----------
Prob(neither fails) = 0.98^2 = 0.9604...
----------------
Prob(one or other fails) = 1 - [0.0004+0.9604] = 0.0392..
-----------------------------------
Cheers,
Stan H.